OpenStudy (anonymous):

Question - 19 : Signals And Systems Using the \(\color{blue}{\text{Convolution Property of Bilateral Laplace Transform}}\), find the response of the system having Impulse Response and Input Signal as: \(a) \large \quad \color{green}{f(t) = e^{\frac{|t|}{4}}}\) \(b) \large \quad \color{red}{h(t) = e^{-3t} \cdot u(t)}\)

OpenStudy (anonymous):

Convolution Property is given as: \[\text{f(t)*h(t) = F(s) H(s)} \quad \quad * \rightarrow Convolution\]

OpenStudy (anonymous):

I just need to check the answer..

OpenStudy (anonymous):

I have calculated it using all my brain..

OpenStudy (dan815):

haha

OpenStudy (dan815):

ok umm soo u want a * a and b * b right

OpenStudy (anonymous):

Nope..

OpenStudy (dan815):

or f(t) * h(t)?

OpenStudy (anonymous):

yep//

OpenStudy (dan815):

ok

OpenStudy (anonymous):

I got it as: \[\large y(t) = f(t)*h(t) = \frac{4}{13}e^{\frac{t}{4}} \cdot u(-t) + \frac{4}{11} e^{\frac{-t}{4}} \cdot u(t) - \frac{8}{143}e^{-3t} \cdot u(t)\]

OpenStudy (anonymous):

I got this weird answer.. ROC is : \(\frac{-1}{4} <\Re\{s\} < \frac{1}{4}\)

OpenStudy (dan815):

ummm i see

OpenStudy (dan815):

what happens if u try to take L {e^-3t u(t)}

OpenStudy (anonymous):

I can give some more details I got while walking through the solution..

OpenStudy (dan815):

isnt there a way to check out answer

OpenStudy (anonymous):

\[H(s) = \frac{1}{s+3},\quad \Re\{s\} > -3\]

OpenStudy (anonymous):

Ha ha ha, For that I have come here.. :P

OpenStudy (anonymous):

\[e^{-at} u(t) \longleftrightarrow \frac{1}{s+a} , \quad \sigma > -a\]

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

you didn't use convolution property ?

hartnn (hartnn):

L(f*g) = FG

OpenStudy (anonymous):

@hartnn I have got that result by using convolution property only.. I just need to check my answer..

OpenStudy (anonymous):

How you got that as your Multiplication Result?

hartnn (hartnn):

this is the property L(f*g) = FG Laplace of convolution is the product of laplaces.

hartnn (hartnn):

you have that right and how did you calculate laplaces of f and h

OpenStudy (anonymous):

I think my question is wrong..

OpenStudy (anonymous):

The bilateral transform of \(e^{\frac{|t|}{4}}\) does not exist I think, @zarkon right?